estion 3 of 5 > Atten Suppose that a simple random sample of size 400 selected from a population, p, has 257 successes. Calculate the has 257 successes. Calculate the margin of error for a 90% confidence interval for the proportion of successes for the population, p. Show the results from each intermediate step performed to calculate the margin of error. First., proportion, , and use the sample proportion to calculate the standard error estimate, SE. Then determine the critical value, 2, using a z-distribution table. Perform interpolation to derive the value of z to three decimal places. Finally, calculate the margin of erro pute the sample r, m. Give all of your answers with precision to three decimal places. SE m- contact us ENG 12:00 AM da) UK 2018-11-17 ^
Answer #1

Ans:

sample proportion,p-257/400 0.6425 0.643

SE=sqrt(0.6425*(1-0.6425)/400)=0.023963\approx 0.024

For 90% confidence interval,z=1.645

Margin of error,m=1.645*0.023963=0.039